A 0.772 gram sample of hydrogen gas has a volume of 856 milliliters at a pressure of 2.79 atm. The temperature of the H2 gas sample is °C.
Answer #1

Molar mass of H2 = 2.016 g/mol

mass(H2)= 0.772 g

use:

number of mol of H2,

n = mass of H2/molar mass of H2

=(0.772 g)/(2.016 g/mol)

= 0.3829 mol

Given:

P = 2.79 atm

V = 856.0 mL

= (856.0/1000) L

= 0.856 L

n = 0.3829 mol

use:

P * V = n*R*T

2.79 atm * 0.856 L = 0.3829 mol* 0.08206 atm.L/mol.K * T

T = 76.0 K

= (76 - 273) oC

= -197 oC

Answer: -197 oC